5.​Calculate the WAAC with the following information:
​​Equity Information​​​​Debt Information
​​10,000 shares​​​​$200,000 in outstanding debt ​​​​​​​​​(face value)
​​$60 per share
​​Beta = 1.2​​​​​Current quote = 100
​​Market risk premium = 12%​​Annual coupon rate = 10%
​​Risk-free rate = 5%​​​​Tax rate = 20%
Answer #1

​WACC=(VE​×Re)+(VD​×Rd×(1−Tc))

WACC formula is the summation of two terms:

\left ( \frac{ E }{ V} \times Re \right )(VE​×Re)

\left ( \frac{ D }{ V} \times Rd \times ( 1 - Tc ) \right )(VD​×Rd×(1−Tc)